# Detecting Shearless Phase-Space Transport Barriers in Global Gyrokinetic Turbulence Simulations with Test Particle Map Models

### v1

@@ 3.2 Planar Map Dynamics and Conditions for Shearless Tori
In this section, we describe how the model dynamical system, which is constructed in the full gyrokinetic phase space, can be reduced to a planar map through an appropriately chosen surface of section, i.e. a Poincaré section. This process is exactly analogous to the usage of surfaces of section to study energetic particle transport under the influence of a single toroidal Alfvén eigenmode (Hsu & Sigmar, 1992; Todo, 2019). Similar to how the safety factor $q$ captures topological information about magnetic field lines through their average winding numbers around the torus, we use the kinetic safety factor $q_{kin}$ (Gobbin et al., 2008) to infer topological information about the test-particle orbits. The presence of non-degenerate minima and maxima in $q_{kin}$ provides a necessary condition for the potential existence of phase-space transport barriers in the form of shearless invariant tori.
@@
which is the ratio of the average toroidal transit frequency $\Omega_{\varphi^{\prime}}$ in the moving frame, i.e. the frequency by which the particle completes a $2\pi$ orbit toroidally in the frame rotating with frequency $\Omega$, to average poloidal return frequency $\Omega_{\theta}$, i.e. the frequency by which the particle takes to return to the midplane.
$q_{kin}$ has a different physical origin for passing and trapped particles due to the topological difference between passing and trapped orbits. For passing particles, the toroidal and poloidal transit times are determined by the parallel transit times of the particle along magnetic field lines as they wind around the torus. Meanwhile for trapped particles, the toroidal transit time is determined by the toroidal precession frequency of the bounce orbits, and the poloidal transit time is given by the bounce time. This definition of $q_{kin}$ is relevant for dynamical systems theory, as it is equal to the rotation number of an orbit under the action of the Poincaré map $f^{\pm}_{\mu,K}$. $q_{kin}$ can be computed using this definition by directly integrating the characteristic equations (2) using an RK4 method for ODEs. Whenever possible for the numerical integration, quantities involving derivatives of fields are expressed analytically in terms of derivatives of interpolating functions of the equilibrium poloidal flux function $\psi(R,Z)$ and flux-surface functions $f(\psi)$.
@@
A second case occurs when $q_{kin}$ has a non-degenerate minimum or maximum at an isolated radial coordinate $P_{\varphi}^{*}$, equivalently $q_{kin}^{\prime}(P_{\varphi}^{*})=0$ and $q_{kin}^{\prime\prime}(P_{\varphi}^{*})\neq 0$. In this case $f^{\pm}_{\mu,K}$ is known as a nontwist map, and the invariant torus labeled by $P_{\varphi}^{*}$ is known as a nontwist or shearless torus. In Appendix A, we note that the definition of shearless torus is independent of the choice of Poincaré section, implying that the nontwist condition is a topological condition applicable to the continuous-time dynamics as well. Despite classical KAM theory being inapplicable, shearless tori are phenomenologically observed to be extremely robust to perturbation, often times being the last remaining invariant torus in otherwise chaotic systems. Physics-focused reviews of shearless tori can be found in del Castillo-Negrete (2000); Morrison (2000); Caldas et al. (2012), and rigorous mathematical justification of their existence has been established with non-twist KAM theorems (Delshams & De La Llave, 2000).
@@
We end this section by remarking that recent work has used singularity theory to classify shearless tori in higher dimension and prove corresponding nontwist KAM theorems about their persistence under perturbation (González-Enríquez et al., 2014). For this work, we primarily focus on planar maps, where much more is known about the phenomenology of shearless tori. However, as pointed out in earlier works del Castillo-Negrete et al. (1996); Rypina et al. (2007b), one way to define shearless tori in integrable systems is through degeneracies in the quadratic part of the Hamiltonian. In higher dimension, these degeneracies can be detected in action-angle coordinates by looking for values of $J$ where the local shear $D\omega(J)$ is singular. Here $D\omega(J)$ is the matrix of partial derivatives $D\omega(J)]_{ij}=\partial_{J_{j}}\omega_{i}$ of the torus frequencies $\omega_{i}=\partial_{J_{i}}H$ with respect to the action coordinates $J_{j}$. Observe that $D\omega(J)$ is equal to the matrix of second derivatives of the Hamiltonian with respect to $J$, so a shearless torus with action $J^{*}$ can be associated with a degenerate critical point of an associated potential function $V(J)=H(J)-\omega(J^{*})\cdot J$, furnishing the link with singularity theory. We will discuss some conjectures relating to these higher-dimensional shearless tori in §5.1.
@@ 3.3 Persistence of Shearless Tori in Model Dynamical System
Now, we consider the full model dynamical system with electrostatic potential $\hat{\phi}=\langle\phi\rangle+\delta\hat{\phi}$ including the ballooning mode (5) extracted from the gyrokinetic simulation. We consider the motion of the deuterium ions, as the modes rotate in the ion diamagnetic direction and interact more strongly with ion trajectories. We take $\mu,K$ to be the same as the passing $\mathcal{E}_{\perp}/\mathcal{E}=1/3$ and trapped $\mathcal{E}_{\perp}/\mathcal{E}=2/3$ cases considered in figure 5. We solve the characteristic equations (2) using an RK4 method for ODEs. Note that previous studies have shown that the gyro-averaging can reduce the level of chaos present in a system (del Castillo-Negrete & Martinell, 2012; Martinell & del Castillo-Negrete, 2013; Da Fonseca et al., 2014; Kryukov et al., 2018), and hence we focus on the case where the gyro-averaging is applied. To implement the gyro-average, we take the eikonal approximation $\mathbf{k}_{\perp}=n\nabla(\varphi-q\eta)$ and multiply the ballooning mode envelope $g_{n}$ by the gyro-average factor $J_{0}(k_{\perp}\rho)$.
To visualize the chaotic transport due to the perturbed electrostatic potential, we plot Poincaré sections in Figure 6. Panels (a,b) shows the Poincaré section for the model gyrokinetic system with the drift wave at the amplitude observed in the simulations for trapped and passing particles corresponding to the $\mathcal{E}_{\perp}/\mathcal{E}=2/3$ and $\mathcal{E}_{\perp}/\mathcal{E}=1/3$ cases. Focusing on panel (a), despite presence of chaos throughout the section, the trapped particles mostly remain in two distinct regions of phase space separated by a radial barrier around $P_{\varphi}/\psi_{lcfs}\approx 0.8$. Plotting $q_{kin}$ in the bottom sub-panel shows that this corresponds to a shearless torus, with a non-degenerate minimum appearing in the rotation number plotted against the radial coordinate. In contrast, panel (b) shows no clear barrier to radial transport for passing particles.
@@
Figure 7: Heatmap showing the tranmissivity $\eta_{t}$ across the shearless transport barrier region varying $v_{\parallel}$ and $v_{\perp}$. Unshaded white regions correspond to $\eta_{t}=0$. The trapped/passing boundary in velocity space is shown with blue lines.
@@
A plot of the transmissivity $\eta_{t}(v_{\perp},v_{\parallel})$ is shown in figure 7. Overplotting the trapped/passing boundary $(v_{\perp}/v_{\parallel})^{2}=B_{max}/B_{min}-1$ reveals that the majority of trapped particle phase space experiences a complete suppression of transport across the shearless transport barrier region, strongly suggestive that shearless invariant tori as identified in figure 6(a,c) are generic across most of trapped particle phase space. We remark there is a band of much larger transmissivity within the trapped portion of phase space near $v_{\perp}/v_{t}\sim 2$, meaning that trapped particles with energies slightly higher than thermal cross the shearless region more easily. This suggests that, in this particular case, the shearless tori primarily affect the density transport channel.
We remark that trapped particles in the shearless region with $v_{\perp}/v_{t}\sim 2$ have nearly stationary toroidal precession in the frame rotating with $\Omega$, indicating strong wave-particle resonance with the model drift wave. More generally, one could expect many features of figure 7 to originate from the structure of particle phase-space resonances; see for example Antonenas et al. (2024) for a detailed study of how $q_{kin}$ determines the structure of the resonances. The reduction in the size of resonances near shearless regions has been proposed as a mechanism by which shearless tori or their remnants can suppress transport (Rypina et al., 2007b). We defer detailed study of these resonances in response to drift-wave perturbations to future work.
@@
We also consider averages of $\eta_{t}$ over a thermal Maxwellian where the integrals are restricted to trapped and passing portions of velocity space respectively. These averages are reported in Table 1. Despite the transmission band for trapped particles around $v_{\perp}/v_{t}\sim 2$, the overall effect of the shearless tori is to reduce the transmissivity by a factor of $2$ in the trapped portions of velocity compared to the passing portions of phase velocity. Trapped particles make up about $75\%$ of thermal particles at the outboard midplane, suggesting that the presence of shearless tori can lead to a significant drop in the overall radial particle transport due to the presence of shearless tori. For steady-state density profiles with no core fueling, the density gradient in the outer core will be determined by the balance between inward particle pinches with outward particle diffusion. Although we have not ruled out a localized increase in the inward particle pinch, the presence of shearless tori suggests a plausible mechanism by which a long-lived increase in ion gyrocenter density gradients can be sustained in the shearless region.
@@ 4 Signatures of Shearless Tori in Gyrokinetic Simulations
Having established the role of shearless tori in suppressing radial transport in the model test particle dynamics, we now turn to compare the result of the model with the fully self-consistent gyrokinetic dynamics simulated by XGC. In this section, we begin in §4.1 by presenting direct evidence of signatures of shearless tori in the gyrokinetic simulations visible in the ion gyrocenter distribution function. In §4.2 we then study the time-dependent behavior of these phase space signatures and demonstrate a link between the arrest of turbulence propagation at the shearless region with reconnection events involving shearless tori in gyrokinetic phase space. These signatures provide evidence of the active role that shearless transport barriers play in regulating the turbulence dynamics, and suggest phenomenological approaches for modeling the interaction between turbulence and shearless regions.
@@ 4.1 Direct Evidence of Shearless Tori in Gyrokinetic Simulations
4.2 Shearless Barrier Reconnection
@@
Having established the presence of shearless phase-space transport barriers, we now turn to the question of the mechanism which stops the radial propagation of avalanches across the shearless region. In this section, we show how avalanches impinging on the shearless tori cause zonal jet reconnection and subsequent eddy detachment, reminiscent of the formation of warm and cold core ring structures in oceanic jets.
@@
Figure 10: (a) Temporal snapshot sequence of ion GC distribution function $F_{i}$. The timestamp is shown in colored text, and a single reconnection event is tracked by the ‘+’ symbols. (b-d) Sequence of panels showing the density skewness along with the results from the persistent homology analysis of blob and hole structures. The ‘+’ symbols indicate the same reconnection event as in the left panel. (e) Image of the Gulf Stream sea surface temperature showing eddy detachment, leading to the formation warm and cold core rings (source: https://earthobservatory.nasa.gov/images/5432/the-gulf-stream)
@@
In Figure 10(a), we show the time evolution of the ion gyrocenter distribution function $F_{i}$ for trapped particles in the $\mathcal{E}_{\perp}/\mathcal{E}=2/3$ case, restricted to $E_{\mu,K}$ as before. For ease of viewing, the data is projected from a poloidal plane to flat coordinates $(\alpha,\psi_{n})$, where $\alpha=\varphi-q\theta$ is the perpendicular field line label. Focusing on the propagation of structures from the edge towards the core, we observe several tongue-like structures that pinch off into blobs that then quickly dissipate. Note that in $E_{\mu,K}$, $F_{i}$ increases radially outward, so phase space structures propagating from the edge correspond to phase space blobs here.
@@
Tracking the evolution of these perturbations in time and radius and overplotting it onto the density skewness in Figure 10(b), we observe that these structures precisely track the radial propagation of the avalanches. Moreover, notice that the blob pinch-off event corresponds with the termination of the density skewness avalanche at the shearless region. We remark that this behavior is highly reminiscent of the eddy detachment observed in oceanic jets, such as the Gulf Stream (The Ring Group, 1981; Olson, 1991). An example of eddy detachment in the Gulf Stream is shown in Figure 10(e). The meandering of the Gulf Stream, associated with north-south intrusions of warmer and colder water from the tropics and arctic respectively, pinches off to form eddy rings with cores of warm or cold water detectable by satellite imaging.
@@
Figure 11: Example 2d function with blobs, both (a) without and (b) with noise applied. An example sublevel set for both functions is shown in orange. (c) Persistence diagram computed by topological data analysis for both original (green crosses) and noisy functions (blue dots), as well as the silhouette (orange curve)
In order to demonstrate that this phase space blob reconnection phenomenon is a robust feature of the data, we apply tools from topological data analysis (TDA) to the ion gyrocenter distribution function $F_{i}$. Specifically, we use sublevel set persistent homology (Mirth et al., 2021) to characterize the presence of robust minima and maxima. An illustration of how sublevel set homology can be used to find robust local minima and maxima of functions is illustrated in figure 11. For a planar function $f(x,y)$, consider the sublevel sets $L^{-}_{s}[f]:=\{(x,y):f(x,y)\leq s\}$. Intuitively, every local maximum of $f(x,y)$ will lead to a range of $s$ values for which $L^{-}_{s}[f]$ will have a “hole”, as illustrated by the orange shaded region in figure 11(a). Mathematically, this means that $L^{-}_{s}[f]$ will have a non-trivial $H_{1}$ homology group. Similarly, every local minimum of $f(x,y)$ will create an isolated connected component which contributes to a non-trivial $H_{0}$ homology group.
TDA provides a quantitative way to address the question of which local maxima and minima of a function are “robust”, as well as providing algorithms for efficiently identifying these features. Sublevel set persistent homology is based on the idea that the topological changes in $L^{-}_{s}[f]$ induced by robust local maxima/minima will persist in the presence of noise. For example, in Figure 11(b) there is still only one “hole” in $L^{-}_{s}[f]$ associated with the most significant of the local maxima, despite the many smaller local maxima/minima created by the noise. The robustness of a given feature is quantified by the persistence P=(DB)/2P=(D-B)/2, where the birth $B$ and death $D$ are the values of $f$ at which the feature first appears and disappears, respectively. Features with stronger persistence are viewed as more robust, as they influence the topology of the sublevel sets for larger ranges of $s$.
@@
The persistence can be plotted against the average birth and death value $\bar{f}=(B+D)/2$ of each feature to create a persistence diagram. The persistence diagrams for both the noisy and non-noisy example $f$ are shown in figure 11(c). Note that persistence diagrams often plot the birth $B$ against the death $D$, and this diagram is simply the 45-degree rotated version of such a diagram. In cases where $f$ has an overall gradient, as is the case for the example function $f(x,y)$, $\bar{f}$ can be used as a proxy for the radial location of the feature in the persistence diagram. Notice in the original function, four blobs are detected around $\bar{f}\approx 0.4$ and two are detected around $\bar{f}\approx 0.8$, which correspond to the two bands of blobs given in the example function. In the noisy version of the function, many more features have been detected, but the major blobs continue to be visible in the persistence diagram. This behavior can be rigorously understood through the stability of persistence diagrams under perturbation (Cohen-Steiner et al., 2007).
@@
Next, in order to give a statistical summary of the persistence diagram into a single curve, we use silhouettes (Chazal et al., 2014), which are computed from these persistence diagrams. This is a curve which is constructed by taking a weighted average of tent functions associated with each feature in the persistence diagram. Silhouettes are used in TDA and machine learning to vectorize the topological information present in a noisy persistence diagram into a single curve. Intuitively, the silhouette is a function which peaks where the persistence diagram is dominated by high-persistence features, and is low where the persistence diagram is dominated by low-persistence features. We use a power weighting of $P^{2}$ to emphasize the high-persistence features. An example silhouette is shown in figure 11(c). Essentially, the key feature of the silhouette is that it captures the two radially distinct bands of blobs as peaks in the silhouette function at the corresponding values of $\bar{f}$.
To put this method into practice, we use the fact that the computational mesh consists of 16 copies of an identical triangular mesh in the poloidal plane, with each mesh corresponding to a different toroidal slice. The ion gyrocenter distribution function $F_{i}$ is available on the nodal points of this mesh. For this analysis we consider just minima and maxima in each isolated 2d poloidal plane. Then, since the computational mesh corresponds to a simplicial complex, we can use existing TDA libraries to compute the sublevel set persistent homology. We use the simplex tree and representations modules of the GUDHI library (Boissonnat & Maria, 2014; Maria, 2025; Dlotko, 2025) to compute the persistence diagrams and silhouettes. In this case, each connected component ($H_{0}$ component) of the sublevel set corresponds to a minimum, while each ‘hole’ ($H_{1}$ component) corresponds to a maximum. An example silhouette computed from the data is shown in Figure 11b, where the silhouette is shown in orange.

### v2

@@ 3.2 Planar Map Dynamics and Conditions for Shearless Tori
In this section, we describe how the model dynamical system, which is constructed in the full gyrokinetic phase space, can be reduced to a planar map through an appropriately chosen surface of section, i.e. a Poincaré section. This process is exactly analogous to the usage of surfaces of section to study energetic particle transport under the influence of a single toroidal Alfvén eigenmode (Hsu and Sigmar, 1992; Todo, 2019). Similar to how the safety factor $q$ captures topological information about magnetic field lines through their average winding numbers around the torus, we use the kinetic safety factor $q_{kin}$ (Gobbin et al., 2008) to infer topological information about the test-particle orbits. The presence of non-degenerate minima and maxima in $q_{kin}$ provides a necessary condition for the potential existence of phase-space transport barriers in the form of shearless invariant tori.
@@
which is the ratio of the average toroidal transit frequency $\Omega_{\varphi^{\prime}}$ in the moving frame, i.e. the frequency by which the particle completes a $2\pi$ orbit toroidally in the frame rotating with frequency $\Omega$, to average poloidal return frequency $\Omega_{\theta}$, i.e. the frequency by which the particle takes to return to the midplane. More generally, $q_{kin}$ is equal to the rotation number of an orbit under the action of the Poincaré map $f^{\pm}_{\mu,K}$, relevant to dynamical systems theory.
$q_{kin}$ has a different physical origin for passing and trapped particles due to the topological difference between passing and trapped orbits. For passing particles, the toroidal and poloidal transit times are determined by the parallel transit times of the particle along magnetic field lines as they wind around the torus. Meanwhile for trapped particles, the toroidal transit time is determined by the toroidal precession frequency of the bounce orbits, and the poloidal transit time is given by the bounce time. In both the integrable and non-integrable cases, $q_{kin}$ can be computed by integrating the characteristic equations (2) using an RK4 method for ODEs. Whenever possible for the numerical integration, quantities involving derivatives of fields are expressed analytically in terms of derivatives of interpolating functions of the equilibrium poloidal flux function $\psi(R,Z)$ and flux-surface functions $f(\psi)$. In the integrable case, the toroidal and poloidal frequencies associated with the orbit can be directly calculated from the trajectories to evaluate (6). In the non-integrable case, $q_{kin}$ can be computed using the weighted Birkhoff average (Sander and Meiss, 2020) to approximate the average rotation number of a particle trajectory under the Poincaré map.
@@
A second case occurs when $q_{kin}$ has a non-degenerate minimum or maximum at an isolated radial coordinate $P_{\varphi}^{*}$, equivalently $q_{kin}^{\prime}(P_{\varphi}^{*})=0$ and $q_{kin}^{\prime\prime}(P_{\varphi}^{*})\neq 0$. In this case $f^{\pm}_{\mu,K}$ is known as a nontwist map, and the invariant torus labeled by $P_{\varphi}^{*}$ is known as a nontwist or shearless torus. In Appendix A, we note that the definition of shearless torus is independent of the choice of Poincaré section, implying that the nontwist condition is a topological condition applicable to the continuous-time dynamics as well. Despite classical KAM theory being inapplicable, shearless tori are phenomenologically observed to be extremely robust to perturbation, often times being the last remaining invariant torus in otherwise chaotic systems. Physics-focused reviews of shearless tori can be found in del-Castillo-Negrete (2000); Morrison (2000); Caldas et al. (2012), and rigorous mathematical justification of their existence has been established with non-twist KAM theorems (Delshams and De La Llave, 2000).
@@
We end this section by remarking that recent work has used singularity theory to classify shearless tori in higher dimension and prove corresponding nontwist KAM theorems about their persistence under perturbation (González-Enríquez et al., 2014). For this work, we primarily focus on planar maps, where much more is known about the phenomenology of shearless tori. However, as pointed out in earlier works del-Castillo-Negrete et al. (1996); Rypina et al. (2007b), one way to define shearless tori in integrable systems is through degeneracies in the quadratic part of the Hamiltonian. In higher dimension, these degeneracies can be detected in action-angle coordinates by looking for values of $J$ where the local shear $D\omega(J)$ is singular. Here $D\omega(J)$ is the matrix of partial derivatives $D\omega(J)]_{ij}=\partial_{J_{j}}\omega_{i}$ of the torus frequencies $\omega_{i}=\partial_{J_{i}}H$ with respect to the action coordinates $J_{j}$. Observe that $D\omega(J)$ is equal to the matrix of second derivatives of the Hamiltonian with respect to $J$, so a shearless torus with action $J^{*}$ can be associated with a degenerate critical point of an associated potential function $V(J)=H(J)-\omega(J^{*})\cdot J$, furnishing the link with singularity theory. We will discuss some conjectures relating to these higher-dimensional shearless tori in §5.1.
@@ 3.3 Persistence of Shearless Tori in Model Dynamical System
Now, we consider the full model dynamical system with electrostatic potential $\hat{\phi}=\langle\phi\rangle+\delta\hat{\phi}$ including the ballooning mode (5) extracted from the gyrokinetic simulation. We consider the motion of the deuterium ions, as the modes rotate in the ion diamagnetic direction and interact more strongly with ion trajectories. We take $\mu,K$ to be the same as the passing $\mathcal{E}_{\perp}/\mathcal{E}=1/3$ and trapped $\mathcal{E}_{\perp}/\mathcal{E}=2/3$ cases considered in figure 5. We solve the characteristic equations (2) using an RK4 method for ODEs. Note that previous studies have shown that the gyro-averaging can reduce the level of chaos present in a system (del-Castillo-Negrete and Martinell, 2012; Martinell and del-Castillo-Negrete, 2013; Da Fonseca et al., 2014; Kryukov et al., 2018), and hence we focus on the case where the gyro-averaging is applied. To implement the gyro-average, we take the eikonal approximation $\mathbf{k}_{\perp}=n\nabla(\varphi-q\eta)$ and multiply the ballooning mode envelope $g_{n}$ by the gyro-average factor $J_{0}(k_{\perp}\rho)$.
To visualize the chaotic transport due to the perturbed electrostatic potential, we plot Poincaré sections in Figure 6. Panels (a,b) shows the Poincaré section for the model gyrokinetic system with the drift wave at the amplitude observed in the simulations for trapped and passing particles corresponding to the $\mathcal{E}_{\perp}/\mathcal{E}=2/3$ and $\mathcal{E}_{\perp}/\mathcal{E}=1/3$ cases. Focusing on panel (a), despite presence of chaos throughout the section, the trapped particles mostly remain in two distinct regions of phase space separated by a radial barrier around $P_{\varphi}/\psi_{lcfs}\approx 0.8$. The bottom sub-panel shows $q_{kin}$ plotted against the average $P_{\varphi}$ for the trajectory, both computed using the weighted Birkhoff average. Although the rotation numbers are perturbed from the integrable case due to the influence of the drift-wave fluctuations, a non-degenerate minimum still appears in the rotation number plotted against the radial coordinate, showing that this barrier corresponds to a shearless torus. In contrast, panel (b) shows no clear barrier to radial transport for passing particles.
@@
Figure 7: (a-b) Heatmaps showing the transmissivity $\eta_{t}$ across the shearless transport barrier region varying $v_{\parallel}$ and $v_{\perp}$, with and without gyroaveraging respectively. Unshaded white regions correspond to $\eta_{t}=0$. The trapped/passing boundary in velocity space is shown with blue lines.
@@
Plots of the transmissivity $\eta_{t}(v_{\perp},v_{\parallel})$ are shown in figure 7, where panel (a) shows the nominal case with gyroaveraging applied to the fields, while panel (b) shows the case with no gyroaveraging applied to the fields. Focusing first on the nominal case in panel (a), overplotting the trapped/passing boundary $(v_{\perp}/v_{\parallel})^{2}=B_{max}/B_{min}-1$ reveals that the majority of trapped particle phase space experiences a complete suppression of transport across the shearless transport barrier region, strongly suggestive that shearless invariant tori as identified in figure 6(a,c) are generic across most of trapped particle phase space. Comparing this to the case with no gyroaveraging in panel (b), the picture is extremely similar, although a close inspection in the trapped region of phase space reveals some regions where the transmissivity has become non-zero.
Both cases show a band of much larger transmissivity within the trapped portion of phase space near $v_{\perp}/v_{t}\sim 2$. Trapped particles in this region of phase space have nearly stationary toroidal precession in the frame rotating with $\Omega$, indicating strong wave-particle resonance with the model drift wave. This indicates that although gyroaveraging may change the critical amplitude for the survival of the shearless torus, proximity to resonance may play a more significant role in determining if the shearless torus survives or not. More generally, one could expect many features of figure 7 to originate from the structure of particle phase-space resonances; see for example Antonenas et al. (2024) for a detailed study of how $q_{kin}$ determines the structure of the resonances. The reduction in the size of resonances near shearless regions has been proposed as a mechanism by which shearless tori or their remnants can suppress transport (Rypina et al., 2007b). We defer detailed study of these resonances in response to drift-wave perturbations to future work.
@@
We also consider averages of $\eta_{t}$ over a thermal Maxwellian where the integrals are restricted to trapped and passing portions of velocity space respectively. These averages are reported in Table 1 for both the nominal case and the case with no gyroaveraging. Despite the transmission band for trapped particles around $v_{\perp}/v_{t}\sim 2$, the overall effect of the shearless tori is to reduce the transmissivity by a factor of $2$ in the trapped portions of phase space compared to the passing portions of phase space. Trapped particles make up about $75\%$ of thermal particles at the outboard midplane, suggesting that the presence of shearless tori can lead to a significant drop in the overall radial particle transport due to the presence of shearless tori. The transmissivity increases slightly without gyroaveraging, although the qualitative picture stays the same.
To connect these results with the observed profiles in the simulations, we posit that the reduction in transmissivity due to shearless tori would likely lead to a drop in the turbulent particle diffusivity $D_{n}$. Meanwhile, the transmission band around $v_{\perp}/v_{t}\sim 2$ suggests that trapped particles with higher energies cross the shearless region more easily, leading to a more pronounced effect in the density transport channel compared to the thermal energy transport channel. For steady-state density profiles with no core fueling, relevant to this scenario, the presence of peaked density profiles requires the existence of an inward flux of particles, also known as a particle pinch (Angioni et al., 2009). The density gradient would be set by the balance between the outward diffusive flux $\sim D_{n}\nabla n$ and the inward pinch $\Gamma_{pinch}$, $\nabla n\sim\Gamma_{pinch}/D$. Although we have not ruled out a localized increase in the inward particle pinch, the presence of shearless tori suggests a plausible mechanism by which a long-lived increase in ion gyrocenter density gradients can be sustained in the shearless region.
@@ 4 Signatures of Shearless Tori in Gyrokinetic Simulations
Having established the role of shearless tori in suppressing radial transport in the model test particle dynamics, we now turn to compare the result of the model with the fully self-consistent gyrokinetic dynamics simulated by XGC. In this section, we begin in §4.1 by presenting direct evidence of signatures of shearless tori in the gyrokinetic simulations visible in the ion gyrocenter distribution function. In §4.2 we then study the time-dependent behavior of these phase space signatures and demonstrate a link between the arrest of turbulence propagation at the shearless region with eddy detachment events involving the shearless tori in gyrokinetic phase space. These signatures provide evidence of the active role that shearless transport barriers play in regulating the turbulence dynamics, and suggest phenomenological approaches for modeling the interaction between turbulence and shearless regions.
@@ 4.1 Direct Evidence of Shearless Tori in Gyrokinetic Simulations
4.2 Shearless Barrier Breakdown via Eddy Detachment
@@
Having established the presence of shearless phase-space transport barriers, we now turn to the question of the mechanism which stops the radial propagation of avalanches across the shearless region. In this section, we show how avalanches impinging on the shearless tori cause eddy detachment, reminiscent of the formation of warm and cold core ring structures in oceanic jets.
@@
Figure 10: (a) Temporal snapshot sequence of ion GC distribution function $F_{i}$. The timestamp is shown in colored text, and a single eddy detachment event is tracked by the ‘+’ symbols. (b-d) Sequence of panels showing the density skewness along with the results from the persistent homology analysis of blob and hole structures. The ‘+’ symbols indicate the same eddy detachment event as in the left panel. (e) Image of the Gulf Stream sea surface temperature showing eddy detachment, leading to the formation warm and cold core rings (source: https://earthobservatory.nasa.gov/images/5432/the-gulf-stream)
@@
In Figure 10(a), we show the time evolution of the ion gyrocenter distribution function $F_{i}$ for trapped particles in the $\mathcal{E}_{\perp}/\mathcal{E}=2/3$ case, restricted to $E_{\mu,K}$ as before. For ease of viewing, the data is projected from a poloidal section to flat coordinates $(\alpha,\psi_{n})$, where $\alpha=\varphi-q\theta$ is the perpendicular field line label. Focusing on the propagation of structures from the edge towards the core, we observe several tongue-like structures that detach into blobs that then quickly dissipate. When restricted to $E_{\mu,K}$, $F_{i}$ increases radially outward, so phase space structures propagating from the edge correspond to phase space blobs here.
@@
Tracking the evolution of one of these perturbations in time and radius and overplotting onto the density skewness in Figure 10(b), we observe that these structures precisely track the radial propagation of the avalanches. Moreover, notice that the blob detachment event corresponds with the termination of the density skewness avalanche at the shearless region. We remark that this behavior is highly reminiscent of the eddy detachment observed in oceanic jets such as the Gulf Stream (The Ring Group, 1981; Olson, 1991). An example of eddy detachment in the Gulf Stream is shown in Figure 10(e). The meandering of the Gulf Stream, associated with north-south intrusions of warmer and colder water from the tropics and arctic respectively, pinches off and detaches to form eddy rings with cores of warm or cold water detectable by satellite imaging.
@@
Figure 11: Example 2d function with blobs, both (a) without and (b) with noise applied. Local maxima are indicated with points in (a), and with crosses ‘+’ in (b). (c) Persistence diagrams computed by sublevel set persistent homology for the original (red points) and noisy functions (red and gray crosses). The silhouette for the noisy case is overplotted (orange curve), indicating two bands of local maxima.
In order to demonstrate that this phase space blob detachment phenomenon is a robust feature of the data, we proceed to study the statistical properties of local maxima and minima in the ion gyrocenter distribution function $F_{i}$. To illustrate the challenge of detecting ‘blobs’ in noisy datasets, we turn to figure 11. Panel (a) shows a function $f$ with two bands of blobs at different values of $x$. In the absence of noise, these blobs can be cleanly identified by finding the local maxima of $f$, marked with red points. However, panel (b) shows the same function $f$ with high-frequency noise added, which may represent noise from particle discreteness or fluctuations from smaller-scale physical processes. Although the two bands of blobs appear to persist, many new local maxima marked with ‘+’ symbols appear due to the presence of noise.
To deal with this challenge, we apply tools from topological data analysis (TDA), namely sublevel set persistent homology (Mirth et al., 2021). Sublevel set persistent homology provides a quantitative way to address the question of which local maxima and minima of a function are “robust”, as well as providing algorithms for efficiently identifying these features. The method is based on computing the homology of sublevel sets, i.e. the number of connected components and holes present in the set of points $(x,y)$ such that $f(x,y)\leq z$ for a given $z$.
To explain how this method quantifies the robustness of local maxima and minima, we give a brief intuitive explanation. If one imagines that $f(x,y)$ measures the height of a landscape and $z$ to be the level of water filling the landscape, then local maxima of $f$ with height greater than $z$ correspond to peaks that form landmasses breaking the surface of the water. For each local maximum, there will exist a range of water heights $f_{birth}\leq z<f_{death}$ where the local maximum will be the tallest peak on its own landmass, where $f_{birth}$ and $f_{death}$ are the values of $f$ for which the landmass is distinct. These values define the persistence $P=(f_{death}-f_{birth})/2$, which measures (half) the range of $z$ values for which the local maximum is the tallest peak on its own distinct landmass. While new local maxima can easily be created by noise, exactly one local maximum will inherit the title of being the tallest peak on a given landmass at a time. Thus, counting high-persistence features gives a way to quantify the number of significant local maxima. A similar intuitive explanation can be constructed for local minima by considering the deepest trench in each distinct body of water.
@@
To use this quantity, the persistence can be plotted against the average birth and death value $\bar{f}=(f_{birth}+f_{death})/2$ of each feature to create a persistence diagram. In the case where $f$ has an overall gradient in $x$, $\bar{f}$ is correlated with the $x$ location of the feature. Note that persistence diagrams also sometimes plot $f_{birth}$ against $f_{death}$, which is simply a 45-degree rotated version of the type of diagram used here. The persistence diagrams for both the noisy and non-noisy example $f$ are shown in figure 11(c). The two bands of blobs are visible as a cluster of high-persistence features detected around $\bar{f}\approx 0.4$ and $\bar{f}\approx 0.8$. Although many more local maxima are present in the noisy function, both persistence diagrams show the same number of high-persistence local maxima corresponding to the two bands of blobs, marked in red. This persistence behavior can be rigorously understood through the stability of persistence diagrams under perturbation (Cohen-Steiner et al., 2007). Physically, $P$ measures (half) the fluctuation amplitude $\delta f$ which is organized into a topologically distinct ‘blob’ of $f(x,y)$.
@@
In order to give a statistical summary of the topological information provided by the persistence diagram into a single curve, we use silhouettes (Chazal et al., 2014). These are constructed by taking a weighted average of tent functions associated with each feature in the persistence diagram. Intuitively, the silhouette is a function which peaks where the persistence diagram is dominated by high-persistence features, and is low where the persistence diagram is dominated by low-persistence features. We use a power weighting of $P^{2}$ to emphasize the high-persistence features. With this weighting, the silhouette $\propto|\delta f|^{3}$, where $\delta f$ is again the fluctuation amplitude associated with topologically distinct blobs or holes. An example silhouette is shown in figure 11(c), showing how the two bands of blobs are captured by peaks in the silhouette curve at the values of $\bar{f}$ corresponding to the radial location of the blobs.